Space Stations

NASA Space Station On-Orbit Status 1 August 2018 – Starboard Solar Array Imagery

By Marc Boucher
Status Report
August 2, 2018
Filed under , ,
NASA Space Station On-Orbit Status 1 August 2018 – Starboard Solar Array Imagery
The SpaceX Dragon resupply ship, on its 15th Commercial Resupply Services mission (CRS-15), is pictured attached to the International Space Station the day after it was captured and installed on the Harmony module. The orbital complex was flying over northern central China near the Mongolian border at the time this photograph was taken. Credit: NASA.

The International Space Station Partners and mission managers polled “go” for tomorrow’s release of the SpaceX Dragon cargo craft for its deorbit and splashdown in the Pacific Ocean to wrap up the CRS-15 mission.
With favorable weather conditions forecast in the splashdown zone, Dragon’s hatch will be closed early Friday around 3 a.m. Central time. Robotic ground controllers will then use the Canadarm2 robotic arm to detach Dragon from the Earth-facing port of the Harmony module around 5:30 a.m. Central time and will maneuver Dragon into the release position.

With Expedition 56 Flight Engineer Serena Auñón-Chancellor of NASA monitoring Dragon’s systems, the ground controllers will release Dragon from Canadarm2 at 11:37 a.m. Central time. After its release, Dragon will conduct a series of “departure” burns to move to a safe distance away from the station.

Several hours later, at 4:23 p.m. Central time, SpaceX flight controllers at Hawthorne, California will command Dragon’s engine to fire for 12 minutes and 53 seconds in the deorbit burn that will enable Dragon to slip out of orbit for its descent back to Earth.

Dragon’s parachute-assisted splashdown is scheduled at 5:17 p.m. Central time, 3:17 p.m. Pacific time, about 410 miles southwest of Long Beach, California. It will take about two days for Dragon to be brought back to port for its cache of cargo and scientific experiments to be unloaded.

On-orbit Status Report

Fluid Shifts: A 55S crewmember performed their Fluid Shifts Dilution Measurements activities today. They collected saliva, blood and urine at multiple points during the day and inserted them into a Minus Eighty Degree Celsius Laboratory Freezer for ISS (MELFI). Fluid Shifts is a joint NASA-Russian experiment divided into Dilution Measurements, Baseline Imaging, and Baseline Imaging with Chibis (Lower Body Negative Pressure). The Fluid Shifts experiment investigates the causes for severe and lasting physical changes to astronaut’s eyes. Because the headward fluid shift is a hypothesized contributor to these changes, reversing this fluid shift with a lower body negative pressure device is being evaluated as a possible intervention. Results from this study may help to develop preventative measures against lasting changes in vision and to prevention of eye damage.

Micro-12: Today the crew terminated specific Micro-12 Group Activation Packs (GAPs) by adding a microbial fixative to preserve the organisms for subsequent ground-based characterization. Micro-12 examines the effects of spaceflight on the physiology of an exoelectrogenic microorganism, Shewanella oneidensis MR-1. Exoelectrogenic microbes can pass electrons through their cell membranes, and so can be used in microbial fuel cells to make electricity from waste organic material. This investigation is expected to provide insight into how microgravity affects biological electron transport systems, and microbial fuel cell use in future space missions.

Aerosol Samplers: Today the crew removed cartridges from Aerosol Samplers (AASs) in Nodes 1 and 3 and then removed and stowed all active and passive samplers. The active battery-powered AASs actively pull in air and collect particles using the principal of thermophoresis. This is accomplished by flowing the sampled air through a large thermal gradient in a narrow channel. During this process, particles are driven to the cold side of the channel, where an electron microscope (EM) grid is held in place by a tiny magnet. When returned to Earth, these grids are placed directly into an EM for analysis.

TangoLab-2: The crew removed three cards and the experiment cubes they contain from the TangoLab-2 facility today. TangoLab-2 is a reconfigurable general research facilities designed for microgravity research and development and pilot manufacturing in the ISS.

Center for the Advancement of Science in Space (CASIS) Protein Crystal Growth (PCG) 11: Today the crew removed the CASIS PCG 11 hardware from a Microgravity Experiment Research Locker / INcubator (MERLIN), deactivated each of the cells on a Handheld High Density Protein Crystal Growth (HDPCG) unit, and then inserted the hardware into a MELFI. CASIS PCG 11 produces acetylcholinesterase crystals, a neurotransmitter enzyme. Crystals grown in microgravity are larger, of higher quality and can be used for a technique called macromolecular neutron crystallography to locate hydrogen atoms in the crystal’s structure. These hydrogen atoms play critical roles in the enzyme’s function and knowing their location clarifies the protein’s structure and function. This research advances the development of better antidotes to fatal organophosphate nerve agents, which act by inhibiting acetylcholinesterase in the nervous system.

NanoRacks Module-9: The crew activated designated mixture tubes for the fifth session of the NanoRacks Module-9 investigation today. The experiments contained in these tubes support a variety of experiments sponsored by the Student Spaceflight Experiments Program (SSEP) National Center for Earth and Space Science Education (NCESSE). The student-designed experiments address challenges of living and working in space. The program is also a key initiative for US science, technology, engineering and math (STEM) education, educating and inspiring the next generation of scientists and engineers to work on the space program. Findings from student experiments may contribute to future experiments that benefit the space program.

Advanced Combustion via Microgravity Experiments (ACME): The crew opened the Combustion Integration Rack (CIR) and replaced one of the ACME controllers. The ACME experiment series being performed in the CIR includes five independent studies of gaseous flames. The primary goals of ACME are to improve fuel efficiency and reduce pollutant production in routine fuel combustion activities on Earth. Its secondary goal is to improve spacecraft fire prevention through innovative research focused on materials flammability.

Lighting Effects: The crew obtained light meter readings in designated ISS locations today. The Lighting Effects investigation studies the impact of the change from fluorescent light bulbs to solid-state light-emitting diodes (LEDs) with adjustable intensity and color. The investigation aims to determine if the new lights can improve crew circadian rhythms, sleep, and cognitive performance.

Biomolecule Extraction and Sequencing Technology (BEST): Today the crew swabbed designated surfaces to collect samples and stowed them in a MELFI. The BEST investigation studies the use of sequencing for the identification of unknown microbial organisms living on the ISS, and how humans, plants and microbes adapt to living on the ISS.

Starboard Solar Array Imagery: Today the crew conducted a photo-survey of the starboard solar array masts and the non-cell side (amber side) of the solar array blankets. Periodic solar array photo surveys are used to quantify any potential damage from Micrometeoroid/Orbital Debris (MMOD) impacts and are a bi-yearly maintenance requirement.

Extravehicular Activity (EVA) Tether Inspection: Today the crew inspected for damage on USOS equipment tethers, and tools with integral tethers, needed during the upcoming Russian EVA 45 on 15 August. Prior to EVA use, soft goods on each Safety Tether, Waist Tether, and D-Ring Extender must be checked for Micrometeoroid/Orbital Debris (MMOD) damage, broken tack stitches, and fraying prior to clearing them for EVA use.

Dragon Cargo Operations: Today the crew finished the last tasks on SpaceX-15 Cargo Message #4, including installation of the LEE Fence Hardware, as they continued to pack and load items slated for return on Dragon. Today’s activities were required before the tomorrow’s activity to install the degraded Space Station Remote Manipulator System (SSRMS) Latching End Effector (LEE) in Dragon. Current weather forecasts are a GO for Dragon departure on 3 August.

Mobile Servicing System (MSS) Operations: Yesterday and overnight, the Robotics Ground Controllers walked the Space Station Remote Manipulator System (SSRMS) off onto Mobile Base System (MBS) Power Data Grapple Fixture (PDGF) 3. The SSRMS was then maneuvered as required to perform the Latching End Effector B (LEE-B) Snare Cables survey before a walk-off from MBS PDGF3 to N2 allowed for controllers to grapple the Dragon vehicle in preparation for Friday’s release.

Today’s Planned Activities:
All activities are complete unless otherwise noted.

Fluid Shifts Saliva Baseline Collection
Fluid Shifts MELFI Insertion Operations
Fluid Shifts Urine Collection
Fluid Shifts MELFI Insertion Operations
Fluid Shifts HRF Centrifuge Frozen Blood Baseline
Fluid Shifts Galley Water Collection Tracer Ingestion Operations
Fluid Shifts HRF Centrifuge Blood Collection Configuration
Fluid Shifts USOS Operations Historical Documentation Photography
Fluid Shifts HRF Centrifuge Blood Collection Spin Conclude
Cargo Transfer to Dragon
Fluid Shifts MELFI Insertion
Cold Stowage Dragon Double Coldbag Pack Overview Review
Troubleshooting C&W panel failure to disable audio tone in DC1.
Environmental Health System Redeploy Day/Night Glasses
Battery Stowage Assembly (BSA) Operation Termination
Micro-12 Sample Activation and Termination
Polar Transfer To Dragon Overview Review
ПХО and DC1 Config for EVA
Space Acceleration Measurement System Screen Cleaning
Lighting Effects Light Meter Setup
Fluid Shifts Urine Collection Mid – Subject
Fluid Shifts MELFI Retrieval And Insertion Operations
Fluid Shifts HRF Centrifuge Frozen Blood 3-Hour Collection Subject
Installation of PLETHYSMOGRAPH unit.
Fluid Shifts HRF Centrifuge Blood Collection Configuration
Fluid Shifts Saliva 3-Hour Collection
Intravehicular Tissue Equivalent Proportional Counter (IV-TEPC) Troubleshooting
Fluid Shifts MELFI Retrieval And Insertion Operations
JEM System Laptop Terminal (SLT) Activate with SLT Linux2
PCG-11 Hardware Deactivation
Fluid Shifts HRF Centrifuge Blood Collection Spin Conclude
Fluid Shifts MELFI Retrieval And Insertion Operations
Insertion of PCG-11 Hardware into MELFI
Fluid Shifts HRF Centrifuge Frozen Blood Collection Conclude And Stow
MERLIN 5 Desiccant Swap
Lighting Effects Light Meter Readings
Photo/TV SAW and MAST Camera Setup
Aerosol Samplers Cartridge Removal
Review EVA Translation Paths and Work Zones using DOUG Application.
Photo/TV STBD SAW and Mast Survey
On Board Training (OBT) Dragon Ground Control Release Self Study
Aerosol Samplers Charger Stow
NanoRacks Module 9 Ops Session 5
Aerosol Passive Samplers Hardware Stowage
Lighting Effects Light Meter Stow
Fluid Shifts Saliva 5-Hour Collection – Subject
Review EVA Translation Paths and Work Zones through ISS Windows.
Fluid Shifts MELFI Retrieval And Insertion Operations
Fluid Shifts Urine Collection End – Subject
Fluid Shifts MELFI Insertion Operations
Photo T/V Dragon HD Video Setup
Dragon Latching End Effector (LEE) Fence Bolting
Radio Frequency Identification Logistics CTB Relocate 1
Combustion Integrated Rack Rack Doors Open
PCG-8 MELFI Insert
Fluid Shifts Urine Collection Stow
Combustion Integrated Rack Front End Cap Open
TangoLab-2 Card Cube Remove
US Extravehicular Activity (EVA) Tether Inspection
Dragon Vestibule Outfitting Kit (VOK) Gather
ACME Controller Replace 1
Removal of two hatch tools from Progress 438 [AO] hatch front, Installation of Progress 438 [AO] docking mechanism
Cold Stowage Dragon Double Coldbag Pack Overview Review
TangoLab-1 Card Cube Remove
BEST Swabbing Operations
Combustion Integrated Rack Front End Cap Close
CALCIUM. Experiment Session 8
BEST MELFI Sample Insertion into a Box Module
Combustion Integrated Rack Doors Close
SpX-15 Dragon Payload Cable Preparation
Colorimetric Water Quality Monitor (CWQMK) Activation and Checkout (ACO)
Photo/TV SAW and Mast photography equipment teardown
Cargo Transfer to Dragon
Polar Express Rack Uninstall, Transfer, Handover And Dragon Install
Polar Express Rack Uninstall, Transfer And Handover
Dragon Cargo Operations Conference

Completed Task List Activities:
SM Printer Swap

Ground Activities:
All activities are complete unless otherwise noted.
Solar Array positioning for survey
EHDC sensor assessment

Three-Day Look Ahead:
Thursday, 08/02: BEAM Ingress, LEE FSE Assembly, LEE install in Dragon, SLS stow in BEAM, BEAM egress, Chemical Garden, Coldbag Packing, AstroPi EPO
Friday, 08/03: CBM Vestibule Ops, Dragon Unberth/Release, BCAT, ACE-M2 Sample, Habitat H2O Fill, EVA tool Gather, mobiPV, Chemical Garden
Saturday, 08/04: Weekly Housekeeping, Crew Off-Duty

QUICK ISS Status – Environmental Control Group:
Component Status
Elektron – On
Vozdukh – Manual
[СКВ] 1 – SM Air Conditioner System (“SKV1”) – Off
[СКВ] 2 – SM Air Conditioner System (“SKV2”) – On
Carbon Dioxide Removal Assembly (CDRA) Lab – Operate
Carbon Dioxide Removal Assembly (CDRA) Node 3 – Standby
Major Constituent Analyzer (MCA) Lab – Operate
Major Constituent Analyzer (MCA) Node 3 – Idle
Oxygen Generation Assembly (OGA) – Process
Urine Processing Assembly (UPA) – Standby
Trace Contaminant Control System (TCCS) Lab – Full Up
Trace Contaminant Control System (TCCS) Node 3 – Off

SpaceRef co-founder, entrepreneur, writer, podcaster, nature lover and deep thinker.